my problem states that the shape of a supporting arch can be modeled by
H(x) = -0.03x^2 + 3x
where h(x) represents the height of the arch and x represents the horizontal distance from one end of the base of the arch in meters. Find the maximum height of the arch.
I need a step by step instruction on how to do these types of problems.

Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

my problem states that the shape of a supporting arch can be modeled by
H(x) = -0.03x^2 + 3x
where h(x) represents the height of the arch and x represents the horizontal distance from one end of the base of the arch in meters. Find the maximum height of the arch.
-------------------
I don't know what level of math you're in, but max and min problems use the 1st derivative of the function.
h(x) = -0.03x^2 + 3x
Set the 1st derivative to zero.
0 = -0.06x + 3
0.06x = 3
x = 50 meters
------------
h(50) = -0.03*2500 + 150
= -75 + 150
h = 75 meters

Answer by stanbon(75887)

H(x) = -0.03x^2 + 3x
where h(x) represents the height of the arch and x represents the horizontal distance from one end of the base of the arch in meters. Find the maximum height of the arch.
------------------
The shape of that equation is a parabola. It has either a maximum
point or a minimum point. Your equation has a maximum.
-----------
That maximum occurs when x = -b/2a = -3/(2*-0.03) = 50
-----------------
So the maximum height is H(50) = -0.03(50)^2 + 3*50 = 75 ft.
========================================================
Cheers,
Stan H.